Grade 12th passGeneral Physics

v1 is the speed of sound in a diatomic gas at 273 C and v2 is the r.m.s speed of its molecules at 273 K,then v1/v2

sri
8 Years agoGrade 12th pass

1 Answer

Saurabh Koranglekar
6 Years ago
V1=273degrees which is equal to 546K
v2=273K
v1/v2=546/273=2